OpenStudy (anonymous):

Solve: 2 1/4 = x - 3 1/2=

OpenStudy (anonymous):

* BTW NO EQUAL SIGN ON THE END OF PROBLEM

OpenStudy (anonymous):

1/2=x - 1.5 x= .5 +1.5 x = 2

OpenStudy (theviper):

\[\frac{21}{4}=x-\frac{31}{2}\]\[=>x=\frac{21}{4}+\frac{31}{2}=?\]

OpenStudy (theviper):

Can u now solve it?

OpenStudy (theviper):

Is that clear????????

OpenStudy (theviper):

Any doubt ask plz:)